OpenStudy (anonymous):

Write the equation of the line that passes by (–1, 4) with a slope of –1 in standard form

OpenStudy (anonymous):

y = -x +3?

OpenStudy (anonymous):

did that in my head...need verification

OpenStudy (anonymous):

a –x + 4y = –1 b x – 4y = –1 c x + y = 3 d x – y = –3 Those are the choices

OpenStudy (anonymous):

i say A

OpenStudy (anonymous):

I hope you're right lol

OpenStudy (anonymous):

just wait a sec...someone will pop in